OpenStudy (anonymous):

Simplify cube root of 7 over fifth root of 7. 7 to the power of 1 over 5 7 to the power of 8 over 15 7 to the power of 5 over 3 7 to the power of 2 over 15 @iGreen. im so confused

OpenStudy (jhannybean):

Can you draw or write your problem out for me?

OpenStudy (anonymous):



OpenStudy (jhannybean):

Alright, now change both the numerator and denominator into fractions.

OpenStudy (anonymous):

do what

OpenStudy (jhannybean):

Remember, \(\large \sqrt[n]{x^m} =x^{m/n}\)

OpenStudy (anonymous):

im really confused about this kinda stuff

OpenStudy (jhannybean):

jus work with me here, you will learn.

OpenStudy (anonymous):

ok

OpenStudy (jhannybean):

Take the numerator, the power ontop of the 7 is 1. The power of the square root is 3. When writing is as a fraction, the power of the root becomes the denominator of the fraction.

OpenStudy (jhannybean):

Now can you write the numerator as a fraction for me?

OpenStudy (anonymous):

uhm so now 7 is 1?

OpenStudy (jhannybean):

7 is not one, just think of the format here: \[\huge \sqrt[\color{red}{n}]{x^\color{blue}{m}} =x^{\color{blue}{m}/\color{red}{n}} \implies \sqrt[\color{red}{3}]{7^\color{blue}{1}}=~?\]

OpenStudy (jhannybean):

Think of your 7 like the x.

OpenStudy (anonymous):

so when i did the math i got the total answer of 5.77377402739is that right

OpenStudy (jhannybean):

There is an easier way to go about finding the answer, just follow my method :P

OpenStudy (anonymous):

but im pretty sure that is correct.

OpenStudy (jhannybean):

Yuor answer does not fit any of your answer choices.

OpenStudy (anonymous):

oh okay

OpenStudy (jhannybean):

So what is the numerator as a fraction? cmon! you got this.

OpenStudy (jhannybean):

If we change both the numerator and denominator into fractions when we can subtract them and it will give us a much more simplified answer.

OpenStudy (jhannybean):

Ok, I will help you with the numerator, but you must write the denominator in the same way.

OpenStudy (anonymous):

okay

OpenStudy (jhannybean):

\[\huge \sqrt[\color{red}{n}]{x^\color{blue}{m}} =x^{\color{blue}{m}/\color{red}{n}} \implies \sqrt[\color{red}{3}]{7^\color{blue}{1}}=7^{\color{blue}{1}/\color{red}{3}}\] Tell me if you understand what I did, because you will write the denominatr in the same way, just like this.

OpenStudy (anonymous):

so the next one will be 7^1/5

OpenStudy (jhannybean):

Yes!! :)

OpenStudy (jhannybean):

So now we have \[\huge \frac{7^{1/3}}{7^{1/5}} = 7^{(1/3) - (1/5)}\]

OpenStudy (jhannybean):

Do you know how to subtract fractions?

OpenStudy (jhannybean):

\[\frac{1}{3}-\frac{1}{5}\]

OpenStudy (anonymous):

find the least common denomiantor then add what you did to the bottom to the top then subtract

OpenStudy (jhannybean):

Ok, and so what will you get? :)

OpenStudy (anonymous):

6/15-3/15

OpenStudy (jhannybean):

not quite 6/15

OpenStudy (anonymous):

5/15

OpenStudy (jhannybean):

Yes :)

OpenStudy (jhannybean):

So (5-3)/15 =?

OpenStudy (anonymous):

2/15

OpenStudy (jhannybean):

Good, :)

OpenStudy (jhannybean):

\[\huge \frac{7^{1/3}}{7^{1/5}} = 7^{(1/3) - (1/5)} = 7^{2/15}\]

OpenStudy (anonymous):

D

OpenStudy (jhannybean):

And which answer choice would that be?

OpenStudy (jhannybean):

Great!